### v1

# Scaling Limit of a Stochastic Clustering Model on $\mathds{R}$

Abstract
We consider an infinite-dimensional stochastic clustering model on $\mathds{R}$. In discrete time, each point of a unit-intensity simple point process moves halfway toward either of its left or right neighbors, chosen uniformly at random. Co-located points are merged into a single point, and the resulting simple point process is rescaled to unit intensity. We show that, when the point processes are shifted so that there is a point at the origin, there is a unique weak limit of these dynamics when the initial point process is renewal. We also conjecture that for the time-reversed process and with an appropriate scaling in space, there is a limiting (random) distribution function on $\mathds{R}$, whose corresponding measure assigns to $\mathds{R}$ a measure corresponding to the gap between consecutive points. Finally, we discuss some relevant research directions.
1 Introduction
Dynamic clustering is often done via numerical optimization [li2021survey]. For large, finite datasets, there are two challenges with this approach. First, large numerical optimization problems are computationally expensive. Second, and perhaps more important, is the question of determining stopping criteria for the clustering algorithm. For finite datasets, continued use of a clustering algorithm often results in all data points being assigned to a single cluster, which is undesirable.
In this work, we address this second question. Specifically, we study stochastic dynamic clustering algorithms on infinite datasets, aiming to determine whether these dynamics possess a stationary measure. If a stationary measure exists, a natural stopping criterion for the clustering dynamics is when one is close to that stationary measure. For large finite datasets, one can also accept the current clustering when the cluster gap distribution is close to that of the stationary measure.
To our knowledge, this is the first work to directly analyze this type of clustering dynamics on an infinite dataset, with the intent of informing behavior on large finite datasets. Indeed, we demonstrate this behavior via simulation for two simple algorithms below. These algorithms are as follows.
  • Algorithm 1: Each point moves halfway towards either its left or right neighbor, chosen uniformly at random. Co-located points are merged. The space is always rescaled so that the point process has unit intensity.
  • Algorithm 2: Each point moves halfway towards either its left or right neighbor, chosen such that conditionally, its movement is mean-zero. Here also, co-located points are merged. The space is always rescaled so that the point process has unit intensity.
Simulation results for the two algorithms are presented in Figure 1. Surprisingly, these two algorithms appear to exhibit different behavior: for Algorithm 1, the limit distribution of the gaps appears independent of the initial distribution, whereas for Algorithm 2, it appears to depend on the initial distribution of the gaps.
(a) Algorithm 1
(b) Algorithm 2
Figure 1: Empirical distribution of point gaps after 20 and 25 iterations, starting from gaps distributed as $\mathrm{Exp}(1)$ and $\mathrm{Unif}(0,2)$.
1.1 Related Work
Multiagent averaging dynamics have been studied extensively from a control perspective through various models with applications in distributed coordination and formation [hendrickx2012convergence, jadbabaie2003coordination], animal flocking [chazelle2014convergence], and opinion dynamics [degroot1974reaching, friedkin1999social, hegselmann2002opinion]. While averaging dynamics were initially used to model consensus in multiagent systems (e.g., robotic rendezvous [bullo2009distributed] or consensus in social networks [degroot1974reaching]), these models were later extended to study disagreement (clustering) in addition to consensus. For instance, in the context of opinion dynamics, Friedkin and Johnsen were the first to take this approach by proposing a model that captures opinion polarization into multiple clusters [friedkin1999social]. This model was later extended by Hegselmann and Krause [hegselmann2002opinion], and is known to belong to a broader class of bounded-confidence dynamics [bernardo2024bounded]. Unfortunately, analyzing the convergence and limit points of such dynamics poses major challenges, in contrast to standard averaging models, whose convergence properties can often be studied using classical tools from Markov chains or graph theory. These difficulties become even more pronounced when additional stochasticity is present in the evolution of the dynamics.
It turns out that averaging dynamics on graphs that depend on agents’ states (e.g., agents’ positions) are closely related to dynamic clustering, especially when the data points are spatially located and evolve over time, and the goal is to cluster the points based on their similarities [aggarwal2003framework, Bhatti2019]. Similarly, in the context of dynamic community detection, one aims to explore how diffusion-based averaging dynamics on time-varying graphs can reveal communities (clusters) that persist over time [delvenne2010stability]. In both such applications, the co-evolution of agents’ states and the network structure introduces a dynamic coupling among the agents’ evolutions, which requires new analytical approaches to characterize the limit points and long-term behavior of these dynamics.
Motivated by these applications, we propose and study a new class of averaging dynamics for spatial data clustering, which can be viewed in the same spirit as the Hegselmann-Krause opinion dynamics over infinitely many agents under a simplified neighborhood structure. To our knowledge, no work in the data clustering literature directly considers an infinite data set. There are, however, some recent works in the probability literature that address problems similar to ours. Baccelli and Khaniha [khaniha2025hierarchical] consider an infinite-dimensional clustering model on $\mathds{R}^{2}$; however, they are unable to establish a scaling limit for their dynamics. Angel, Ray, and Spinka [angel2023tale] study a model on $\mathds{R}$ that behaves similarly to clustering, although it is not strictly a clustering model. They leave it as an open problem to determine whether a scaling limit exists for their dynamics. Even identifying the correct scaling is a challenging problem for either of these two papers.
1.2 Contributions and Organization of this Paper
In this paper, we prove that there is a scaling limit for Algorithm 1. That is, (under mild assumptions on the initial data), repeated application of Algorithm 1 converges to a unique limit which is independent of the initial data.
This paper is organized as follows. In Section 2, we describe the model and main results. In Section 3, we prove the main results. In Section 4, we state a conjecture which generalizes our main results. In Section 5, we conclude and discuss future work.
2 Model and Main Results
2.1 Point Process Model
We consider the following model. Let $\mathbf{\Xi}$ be the space of unit-intensity stationary point processes on $\mathds{R}$. Note that $\Xi\in\mathbf{\Xi}$ need not have a point at the origin.
Let $\overrightarrow{\Xi}^{(0)}\in\mathbf{\Xi}$. We define first a point process $\hat{\Xi}^{(t+1)}$ in terms of $\overrightarrow{\Xi}^{(t)}$ as follows. Each point moves halfway towards one of its neighbors, chosen uniformly at random and independently of other points’ movements. If two points move to the same location, they merge so that $\hat{\Xi}^{(t+1)}$ is a simple stationary point process with intensity $\frac{3}{4}$. Then,
$\overrightarrow{\Xi}^{(t+1)}=\frac{3}{4}\times\hat{\Xi}^{(t+1)}$
is obtained by re-scaling $\hat{\Xi}^{(t+1)}$ to unit intensity. Here, we denote the re-scaling to unit intensity by left-multiplying a point process by its intensity: that is, if $\Xi=(\Xi_{i})_{i\in\mathds{Z}}$, then $c\times\Xi:=(c\Xi_{i})_{i\in\mathds{Z}}$, for any constant $c>0$.
Figure 2 shows our dynamics, without the re-scaling to unit-intensity:
Figure 2: A simulation of the point process model, without space re-scaling.
2.2 Gap Sequence Model
We consider a second formulation of the model, in which the dynamics can be considered as a product of certain $\mathds{Z}$–indexed random matrices. In this formulation of the model, rather than track the point locations, we track the gaps between points. The original point process model can be recovered by also tracking a real-valued quantity representing the location of the 0–indexed point.
Indeed, let $\overrightarrow{\Gamma}^{(t)}\in\mathds{R}^{\mathds{Z}}$ be the gap sequence associated to $\overrightarrow{\Xi}^{(t)}$; that is,
$\overrightarrow{\Gamma}^{(t)}_{i}:=\overrightarrow{\Xi}^{(t)}_{i+1}-\overrightarrow{\Xi}^{(t)}_{i}$.
Notice that the components of $\overrightarrow{\Gamma}^{(t)}$ are identically distributed, but generally not independent. In what follows, we treat $\overrightarrow{\Gamma}^{(t)}$ as a $\mathds{Z}$–indexed column vector.
We define two random linear operators, $\mathcal{A}^{(t)}$ and $\mathcal{F}^{(t)}$ as follows, which refer to as averaging and folding, respectively. One iteration of the dynamics is expressed by the dynamics
$\overrightarrow{\Gamma}^{(t+1)}=\mathcal{F}^{(t)}\mathcal{A}^{(t)}\overrightarrow{\Gamma}^{(t)},$
where averaging corresponds to the point movements, and folding corresponds to merging of co-located points. We express these operators as $\mathds{Z}^{2}$–indexed random matrices. To do this, we use the notation $e_{i}$ to represent the $\mathds{Z}$–indexed standard basis vector equal to $1$ in the $i$-th component and equal to 0 in all other components. We also use the notation $\mathcal{A}^{(t)}_{i},\mathcal{F}^{(t)}_{i}$ to denote the $i$-th rows of $\mathcal{A}^{(t)},\mathcal{F}^{(t)}$ respectively, where $i\in\mathds{Z}$. Note that, we have for any $t\geqslant 1$,
$\overrightarrow{\Gamma}^{(t)}=\mathcal{F}^{(t-1)}\mathcal{A}^{(t-1)}\cdots\mathcal{F}^{(2)}\mathcal{A}^{(2)}\mathcal{F}^{(1)}\mathcal{A}^{(1)}\overrightarrow{\Gamma}^{(0)},$ (2.1)
which implies that for any $i\in\mathds{Z}$, the element $\overrightarrow{\Gamma}^{(t)}_{i}$ can be written as a non-negative random linear combination of the entries in $\overrightarrow{\Gamma}^{(0)}$.
Averaging: The averaging operator is defined row-wise as follows, which is clear from the dynamics. For all $i\in\mathds{Z}$,
$\mathcal{A}^{(t)}_{i}:=\begin{cases}\frac{1}{2}(e_{i}+e_{i+1})&\text{ with probability }1/2\\ \frac{1}{2}(e_{i}+e_{i-1})&\text{ with probability }1/2.\end{cases}$
Here, the randomness in each row is independent.
Folding: Observe that point merges at time $t$ occur according to a stationary $\mathds{Z}$–valued renewal process $\tau^{(t)}$, and that for $t\neq s$, $\tau^{(t)}$ and $\tau^{(s)}$ are i.i.d.. It is easy to see that the inter-renewal distribution of $\tau^{(t)}$ is the sum of two independent $\operatorname{Geom}(1/2)$ random variables. From here, it follows that the rows of the folding matrix $\mathcal{F}^{(t)}$ can be expressed as follows. For a process $\tau\subseteq\mathds{Z}$, let $N_{\tau}(\cdot)$ denote the counting distribution of $\tau$, where
$N_{\tau}(i)=\left|\tau\cap[0,i]\right|-\left|\tau\cap[-i,0)\right|\text{ for }i\in\mathds{Z}.$
We have, with $\tau=\tau^{(t)}$,
$\mathcal{F}^{(t)}_{i}:=\begin{cases}\frac{3}{4}(e_{i+N_{\tau}(i)-1}+e_{i+N_{\tau}(i)})&\text{ for }i\in\tau^{(t)}\cap[0,\infty)\\ \frac{3}{4}(e_{i+N_{\tau}(i)}+e_{i+N_{\tau}(i)+1})&\text{ for }i\in\tau^{(t)}\cap(-\infty,0)\\ \frac{3}{4}e_{i+N_{\tau}(i)}&\text{ for }i\not\in\tau^{(t)}.\end{cases}$
At the indices of the mergings, we need to correspondingly add the two merged gaps; otherwise, we keep the original gap. In order to do this, we also need to keep track of the number of mergings over space.
2.3 Time Reversal of Gap Sequence Model
The forward-time gap sequence model consists of averaging and then folding. Thus, the time reversal consists of “un–folding,” and then “un–averaging.” The reverse-time process is as follows.
Recall that in the forward dynamics, $1/4$ of the points are lost due to merging. In the time-reversal, $1/3$ of points should thus split into two new points. Indeed, in the forward process, the inter-duration between points which merge with their left neighbor is given by the sum of two independent $\operatorname{Geom}(1/2)$ random variables. Thus, in the time-reversed process, points “un-merge” with inter-duration $\operatorname{Geom}(1/2)+\operatorname{Geom}(1/2)-1$, where the geometrics are again independent. Thus, let $\rho^{(t)}$ be i.i.d. renewal processes with gap distribution
$R\sim\operatorname{Geom}(1/2)+\operatorname{Geom}(1/2)-1$
where the geometric random variables are independent and take the minimum value $1$. We have $\operatorname{\mathds{E}}R=2+2-1=3$. The requisite coupling between the merged indices and the point processes $\rho^{(t)}$ used in the un-merging is obvious and we omit it for the sake of brevity; see instead Figure 3.
Figure 3: Figure of forward and reverse dynamics.
Let $\mathcal{G}$ denote the space of nonnegative integer-valued sequences indexed by $\mathds{Z}$. The reverse-time dynamics can be expressed in terms of a $\mathcal{G}$–valued Markov process $\eta^{(t)}$. The interpretation of $\eta^{(t)}$ is that each element $\eta^{(t)}_{i}$ is a weight, and the gaps at time $t$ can be constructed by summing and re-scaling the appropriate subset of weights. The exposition is cleaner with integer-valued weights; since averaging induces a factor of $\frac{1}{2}$, we will need to re-scale the integer-valued weights by $\frac{3}{8}$, instead of $\frac{3}{4}$, to compensate for this. The process $\overleftarrow{\eta}^{(t)}\in\mathds{Z}_{+}^{\mathds{Z}}$ evolves as follows.
  • For indices $i\not\in\rho^{(t+1)}$, replace the element $\overleftarrow{\eta}^{(t)}_{i}$ by the tuple $(\overleftarrow{\eta}^{(t)}_{i},\overleftarrow{\eta}^{(t)}_{i})$. This corresponds to “un-averaging.”
  • For indices $i\in\rho^{(t+1)}$, replace the element $\overleftarrow{\eta}^{(t)}_{i}$ by the tuple $(\overleftarrow{\eta}^{(t)}_{i},2\overleftarrow{\eta}^{(t)}_{i},\overleftarrow{\eta}^{(t)}_{i})$. This corresponds to “un-folding” and then “un-averaging.”
  • Finally, construct a new sequence $\overleftarrow{\eta}^{(t+1)}$ by adding the right-most element of the $i$-th tuple with the left-most element of the $(i+1)$-th tuple and removing the parentheses.
Here, this summation corresponds to the fact that the same points contribute to two different gaps after one iteration of the time-reversed dynamics. The sequence $\overleftarrow{\eta}^{(t+1)}$ is indexed so that the 0–indexed point is the sum of the left-most element of the 0–th tuple and the right-most element of the (1)(-1)–st tuple. For example, in Figure 3, the un-mergings correspond to the entire gap which gets deleted by the merging, but also half of each of the preceding and following gaps.
2.4 Main Results
Theorem 2.1.
Let $\overrightarrow{\Xi}^{(0)}$ be a renewal process with finite inter-renewal variance. The following holds:
$\overrightarrow{\Xi}^{(t)}\xrightarrow{\text{f.d.d.}}\overrightarrow{\Xi}^{(\infty)},$ (2.2)
where the limit is independent of $\overrightarrow{\Xi}^{(0)}$. If $\Theta$ is the Palm shift, then
$\Theta\overrightarrow{\Xi}^{(t)}\Rightarrow\Theta\overrightarrow{\Xi}^{(\infty)}.$
We do not actually know the distribution of the Palm shift. The limit is not renewal. Thus, it is difficult to identify and is left as an open problem. The following corollary is a consequence of the proof technique of Theorem 2.1
Corollary 2.2.
The time-reversed process $\left({3}/{8}\right)^{t}\eta^{(t)}$ is the stochastic dual of the gap sequence model, with respect to the inner product
$h(\Gamma^{(s)},\left({3}/{8}\right)^{t}\eta^{(t)}):=\left\langle\Gamma^{(s)},\left({3}/{8}\right)^{t}\eta^{(t)}\right\rangle=\sum_{i\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\Gamma^{(s)}_{i}\eta^{(t)}_{i}$.
Theorem 2.3.
Let $\overrightarrow{G}^{(t)}$ be the number of points of $\overrightarrow{\Xi}^{(t)}$ which have merged with the 0–indexed point of $\Xi^{(0)}$. Then
$\left(\frac{3}{4}\right)^{t}\overrightarrow{G}^{(t)}\xrightarrow{L^{p}}\overrightarrow{G}^{(\infty)},$
for $2\leqslant p<\infty$.
The following corollary follows by coupling the objects in the Proofs of Theorems 2.1 and 2.3; but it is not a direct consequence of the theorem statements themselves.
Corollary 2.4.
The joint weak convergence holds:
$\left(\Theta\overrightarrow{\Xi}^{(t)},\left(\frac{3}{4}\right)^{t}\overrightarrow{G}^{(t)}\right)\Rightarrow\left(\Theta\overrightarrow{\Xi}^{(\infty)},\overrightarrow{G}^{(\infty)},\right).$
3 Proofs of Main Results
3.1 Proof of Theorem 2.1
Recall that $\overleftarrow{\eta}^{(t)}$ is a sequence of weights, from which we will construct (after a re-scaling by $\left(\frac{3}{8}\right)^{t}$ the gap sequence of the limiting point process. Note that $\overleftarrow{\eta}^{(t)}$ is supported on a subset of $\{-2^{t},-2^{t}+1,\ldots,2^{t}\}$, so that
$|\overleftarrow{\eta}^{(t)}|:=\sum_{i=-\infty}^{\infty}{\overleftarrow{\eta}^{(t)}_{i}}$$\infty$.
Define for $t\geqslant 0$ the process
$\overleftarrow{M}^{(t)}:=\left(\frac{3}{8}\right)^{t}\sum_{i=-\infty}^{\infty}{\overleftarrow{\eta}^{(t)}_{i}}$$\infty$.
We can express the evolution of $\overleftarrow{M}^{(t)}$ as follows:
$\overleftarrow{M}^{(t+1)}$ $=\left(\frac{3}{8}\right)^{t+1}\left(2|\overleftarrow{\eta}^{(t)}|+2\sum_{i\in\rho^{(t+1)}}\overleftarrow{\eta}^{(t)}_{i}\right)$
$=\frac{3}{4}\left(\overleftarrow{M}^{(t)}+\left(\frac{3}{8}\right)^{t}\sum_{i\in\mathrm{supp}\overleftarrow{\eta}^{(t)}}\overleftarrow{\eta}^{(t)}_{i}\mathds{1}_{i\in\rho^{(t+1)}}\right)$
$=\overleftarrow{M}^{(t)}+\frac{3}{4}\sum_{i\in\mathrm{supp}\overleftarrow{\eta}^{(t)}}\overleftarrow{X}^{(t)}_{i}\left(\mathds{1}_{i\in\rho^{(t+1)}}-\frac{1}{3}\right)=:\overleftarrow{M}^{(t)}+\overleftarrow{N}^{(t+1)}.$
Here, $\overleftarrow{X}^{(t)}:=\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)}$. Notice that $\overleftarrow{N}^{(t+1)}$ has conditional mean 0. It follows that the process $\overleftarrow{M}^{(t)}$ is a unit-mean positive martingale; hence it converges to an a.s. limit $\overleftarrow{M}^{(\infty)}$.
Notice that since $\max_{i\in\mathds{Z}}\overleftarrow{\eta}^{(t)}_{i}\leqslant 2^{t}$, we also have that $\max_{i\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)}_{i}\leqslant\left(\frac{3}{4}\right)^{t}$. Equivalently, we have $\max_{i\in\mathds{Z}}X_{i}^{(t)}\leqslant\left(\frac{3}{4}\right)^{t}$. Thus,
$\sum_{i\in\mathds{Z}}\left(\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta^{(t)}_{i}}\right)^{2}\leqslant\left(\frac{3}{4}\right)^{t}\overleftarrow{M}^{(t)},$
which converges a.s. to 0 at an exponential rate due to the a.s. convergence of $\overleftarrow{M}^{(t)}$ to $\overleftarrow{M}^{(\infty)}$.
We will now show the $L^{2}$ convergence of $\overleftarrow{M}^{(t)}$ to $\overleftarrow{M}^{(\infty)}$. Indeed, notice that
$\operatorname{\mathds{E}}\left[\left(\overleftarrow{M}^{(t+1)}\right)^{2}\ \biggl|\ \mathcal{F}^{(t)}\right]$ $=\left(\overleftarrow{M}^{(t)}\right)^{2}+\operatorname{\mathds{E}}\left[\left(\overleftarrow{N}^{(t+1)}\right)^{2}\ \biggl|\ \mathcal{F}^{(t)}\right],$
since $\overleftarrow{N}^{(t+1)}$ is conditionally 0 mean. Substituting, we have:
$\operatorname{\mathds{E}}\left[\left(\overleftarrow{N}^{(t+1)}\right)^{2}\ \biggl|\ \mathcal{F}^{(t)}\right]$ $\leqslant\left(\frac{3}{4}\right)^{2}\operatorname{\mathds{E}}\sum_{i\in\mathrm{supp}{\overleftarrow{\eta}^{(t)}}}\sum_{j\in\mathrm{supp}{\overleftarrow{\eta}^{(t)}}}\overleftarrow{X}_{i}^{(t)}\overleftarrow{X}_{j}^{(t)}\left(\mathds{1}_{i\in\rho^{(t+1)}}-\frac{1}{3}\right)\left(\mathds{1}_{j\in\rho^{(t+1)}}-\frac{1}{3}\right)$
$\leqslant\sum_{i\in\mathrm{supp}{\overleftarrow{\eta}^{(t)}}}\sum_{j\in\mathrm{supp}{\overleftarrow{\eta}^{(t)}}}\left(\frac{3}{4}\right)^{2t}\operatorname{Cov}(\mathds{1}_{i\in\rho^{(t+1)}},\mathds{1}_{j\in\rho^{(t+1)}})$
$\leqslant\left(\frac{3}{4}\right)^{2t}\sum_{i\geqslant 1}i\cdot\left|\operatorname{Cov}\left(\mathds{1}_{0\in\rho^{(t+1)}},\mathds{1}_{i\in\rho^{(t+1)}}\right)\right|\to 0.$
Here, the convergence occurs using the fact that the renewal distribution of $\rho$ has an exponential tail and thus, $\operatorname{Cov}(\mathds{1}_{i\in\rho^{(t)}},\mathds{1}_{j\in\rho^{(t)}})$ decays exponentially in |ji||j-i|. It follows that
$\sup_{t\geqslant 0}\operatorname{\mathds{E}}\left(\overleftarrow{M}^{(t)}\right)^{2}<\infty,$
which also establishes the $L^{2}$ convergence of $\overleftarrow{M}^{(t)}$ to a finite unit-mean limit.
Figure 4: Relationship between the weight sequences $\eta^{(t)}$ and the initial gap sequence $\Gamma^{(0)}$. The orange dashed line, for example, is given by $\frac{1}{2}\overleftarrow{\Gamma}^{(0)}_{-1}+\overleftarrow{\Gamma}^{(0)}_{0}+\frac{1}{2}\overleftarrow{\Gamma}^{(0)}_{1}$; the blue dashed line is given by $\frac{1}{2}\overleftarrow{\Gamma}^{(0)}_{0}+\overleftarrow{\Gamma}^{(0)}_{1}+\frac{1}{2}\overleftarrow{\Gamma}^{(0)}_{2}$.
Now, if the gap distribution of $\Xi_{0}$ has finite variance, we have the following (see Figure 4):
$\left\langle\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)},\overleftarrow{\Gamma}^{(0)}\right\rangle$ $:=\sum_{i\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)}_{i}\overleftarrow{\Gamma}^{(0)}_{i}=\overleftarrow{M}^{(t)}+\sum_{i\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)}_{i}(\overleftarrow{\Gamma}^{(0)}_{i}-1)=:\overleftarrow{M}^{(t)}+\overleftarrow{L}^{(t)}.$
Obviously, we have $\operatorname{\mathds{E}}\overleftarrow{L}^{(t)}=0$ for all $t$, since $\operatorname{\mathds{E}}\Gamma^{(0)}_{i}=1$ for all $i\in\mathds{Z}$. Noting that $(\overleftarrow{\Gamma}_{i}^{(0)})_{i\in\mathds{Z}}$ are i.i.d. with variance $\sigma^{2}$ gaps, we have
$\operatorname{\mathds{E}}\left[\left(\overleftarrow{L}^{(t)}\right)^{2}\ \biggl|\ \mathcal{F}^{(t)}\right]$ $=\sum_{i\in\mathds{Z}}\sum_{j\in\mathds{Z}}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta_{i}}^{(t)}\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}_{j}^{(t)}\cdot\operatorname{Cov}(\overleftarrow{\Gamma}_{i}^{(0)},\overleftarrow{\Gamma}_{j}^{(0)})$
$=\sum_{i\in\mathds{Z}}\left(\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta_{i}}^{(t)}\right)^{2}\sigma^{2}\leqslant\left(\frac{3}{4}\right)^{t}\overleftarrow{M}^{(t)}\sigma^{2}\xrightarrow{a.s.}0;$
and thus $\overleftarrow{L}^{(t)}\to 0$ in $L^{2}$. It follows that $\langle\left(\frac{3}{8}\right)^{t}\overleftarrow{\eta}^{(t)},\overleftarrow{\Gamma}^{(0)}\rangle\to\overleftarrow{M}^{(\infty)}$ in $L^{2}$. This establishes the single-dimensional distribution convergence since the choice of $\overleftarrow{\eta}^{(0)}=e_{0}$ is without loss of generality.
Finally, we extend to finite-dimensional distribution convergence. Indeed, for $\overleftarrow{\eta_{0}}=\sum_{i\in F}e_{i}$, where $F$ is finite, we can construct a finite-dimensional martingale driven by the same renewal processes $\rho^{(t)}$, each of which converges a.s. and in $L^{2}$. This last statement follows from shift-invariance and using the previous argument for $\overleftarrow{\eta}_{0}=e_{0}$. It is easily seen that the expectation of all linear functionals of the finite-dimensional process also converge; this establishes the claimed finite-dimensional convergence.
3.2 Proof of Theorem 2.3
The proof of this result is almost identical to that of Theorem 2.1. When defining the reverse-time process, instead of using the tuples $(\overleftarrow{\eta_{i}}^{(t)},\overleftarrow{\eta_{i}}^{(t)})$ and $(\overleftarrow{\eta_{i}}^{(t)},2\overleftarrow{\eta_{i}}^{(t)},\overleftarrow{\eta_{i}}^{(t)})$, we instead use the tuples $(\overleftarrow{\eta_{i}}^{(t)},\overleftarrow{\eta_{i}}^{(t)})$ and $(0,2\overleftarrow{\eta_{i}}^{(t)},2\overleftarrow{\eta_{i}}^{(t)})$.
3.3 Proof of Corollary 2.2
This result follows by checking that this is true when $\eta^{(0)}=e_{i}$ for $i\in\mathds{Z}$ and then appealing to linearity of expectation. Recall that for $h(\cdot,\cdot)$ defined as the inner product in the corollary statement, the Markov process $\Gamma^{(\cdot)}$ and $\eta^{(\cdot)}$ are dual if for all $t,\Gamma^{(0)},\eta^{(0)}$, we have:
$\operatorname{\mathds{E}}_{\Gamma^{(0)}}h(\eta^{(0)},\Gamma^{(t)})=\operatorname{\mathds{E}}_{\eta^{(0)}}h(\eta^{(t)},\Gamma^{(0)}).$
For $\overleftarrow{\eta}^{(0)}=e_{i}$, the result is shown in the proof of Theorem 2.1.
3.4 Proof of Corollary 2.4
This result follows by taking the proofs of Theorems 2.1 and 2.3, and coupling the point processes used to determine the un-merging indices.
4 A Further Conjecture
Let $\mathcal{G}_{+}$ denote the space of non-negative integer-valued sequences indexed by $\mathds{Z}_{+}$.
Conjecture 1.
[…]
Then $\overleftarrow{F}^{(t)}\Rightarrow\overleftarrow{F}^{(\infty)}$, where the (random) mass of $\overleftarrow{F}^{(\infty)}$ is distributed as $\overrightarrow{\Xi}^{(\infty)}_{1}-\overrightarrow{\Xi}^{(\infty)}_{0}$, and the length of the (random) support of $\overleftarrow{F}^{(\infty)}$ is distributed as $\overrightarrow{G}^{(\infty)}_{1}$.
We explicitly write out a dynamics for $\overleftarrow{F}^{(t)}$ in the following lemma.
Lemma 4.1.
We have for all $t\geqslant 0,x\in\mathds{R}$,
$\overleftarrow{F}^{(t)}(x)+\left(\frac{3}{4}\right)^{(t+1)}\sum_{n\leqslant\left(\frac{4}{3}\right)^{t}x}{\left(\frac{4}{3}\right)^{t}\left[\overleftarrow{F}^{(t)}\left(\left(\frac{3}{4}\right)^{t}n\right)-\overleftarrow{F}^{(t)}\left(\left(\frac{3}{4}\right)^{t}(n-1)\right)\right]\left(\mathds{1}_{n\in\rho^{(t+1)}}-\frac{1}{3}\right)}$
$=\overleftarrow{F}^{(t+1)}\left(x+\left(\frac{3}{4}\right)^{t+1}\left(\left|\rho^{(t+1)}\cap\left[0,\left(\frac{4}{3}\right)^{t}x\right]\right|-\frac{1}{3}\left(\frac{4}{3}\right)^{t}x\right)\right).$
For notational simplicity, we denote these dynamics as:
$\overleftarrow{F}^{(t)}(x)+\alpha^{(t+1)}_{x}=\overleftarrow{F}^{(t+1)}(x+\beta^{(t+1)}_{x})$
where
$\alpha^{(t+1)}_{x}:=\left(\frac{3}{4}\right)^{(t+1)}\sum_{n\leqslant\left(\frac{4}{3}\right)^{t}x}{\left(\frac{4}{3}\right)^{t}\left[\overleftarrow{F}^{(t)}\left(\left(\frac{3}{4}\right)^{t}n\right)-\overleftarrow{F}^{(t)}\left(\left(\frac{3}{4}\right)^{t}(n-1)\right)\right]\left(\mathds{1}_{n\in\rho^{(t+1)}}-\frac{1}{3}\right)}$
and
$\beta^{(t+1)}_{x}:=\left(\frac{3}{4}\right)^{t+1}\left(\left|\rho^{(t+1)}\cap\left[0,\left(\frac{4}{3}\right)^{t}x\right]\right|-\frac{1}{3}\left(\frac{4}{3}\right)^{t}x\right).$
Proof.
Fix $t$ and consider some fixed integer $n\in\left[0,\left(4/3\right)^{t}\right]$. There is
$\overleftarrow{Z}_{n}^{(t)}:=\overleftarrow{F}^{(t)}\left(\left(3/4\right)^{t}n\right)-\overleftarrow{F}^{(t)}\left(\left(3/4\right)^{t}(n-1)\right)$
mass at $n$; and by re-scaling by a factor of $(\frac{4}{3})^{t}$, any new mass has the same units over time. If there is an un-merging at the $n$-th index, then an additional $\overleftarrow{Z}^{(t)}_{n}$ is created, and no new mass is created otherwise. However, any new mass is also shrunken by a factor of $\frac{3}{4}$ at time-step $t+1$. Thus, the contribution of the mass at index $n$ at time $t+1$ is given by $\left(\frac{3}{4}\right)^{t+1}\overleftarrow{Z}^{(t)}_{n}(\mathds{1}_{n\in\overleftarrow{\rho}^{(t+1)}}-\frac{1}{3})$. The total additional mass at time $(t+1)$ contributed from indices contributing to $\overleftarrow{F}^{(t)}(x)$ is then given by summing over $n$.
By a similar re-scaling argument, there are $|\overleftarrow{\rho}^{t+1}\cap[0,(\frac{4}{3})^{t}x]|$ un-mergings in this interval; and once again, those are scaled by an additional factor of $\frac{3}{4}$ at time $t+1$.
The right-hand side accounts for the contribution of new mass within the interval $0,(4/3)^{t}x]$ at time $t$ to the distribution function $F^{(t+1)}(\cdot)$. We need to account for the extra mass created at time $t$; this results in a change to the argument of $F^{(t+1)}(\cdot)$. To do this, we count the number of gaps that occur in the interval $0,(4/3)^{t}x]$. This increases by one at each point of $\overleftarrow{\rho}^{t+1}\cap[0,(\frac{4}{3})^{t}x]$, but again any new mass is shrunken by a factor of $\frac{3}{4}$ at time step $t+1$. Thus, the total number of gaps, after re-scaling, is given by $x+\beta^{(t+1)}_{x}$. The result follows. $\blacksquare$
The conjectured convergence of the measures $\overleftarrow{F}^{(t)}$ should follow from the pointwise convergence of their Laplace transforms. Indeed, define for $s\geqslant 0$,
$G^{(t)}(s):=\int_{\mathds{R}}e^{-sx}dF^{(t)}(x).$
Using the dynamics for $F$, we get:
$G^{(t+1)}(s)$ $=\int_{\mathds{R}}e^{-sx}dF^{(t+1)}(x)$
$=\int_{\mathds{R}}e^{-s(x+\beta^{(t+1)}_{x})}dF^{(t+1)}(x+\beta^{(t+1)}_{x})$
$=\int_{\mathds{R}}e^{-sx-s\beta^{(t+1)}_{x}}dF^{(t)}(x)+\int_{\mathds{R}}e^{-sx-s\beta^{(t+1)}_{x}}d\alpha^{(t+1)}_{x}$
$=:H_{\beta}^{(t+1)}(s)+H_{\alpha}^{(t+1)}(s),$
where we treat $\alpha^{(t)}(x)$ as a signed measure. Here, the second equality follows since the mapping $x\mapsto x+\beta^{(t+1)}_{x}$ is monotone increasing in $x$: we have $x+\beta^{(t+1)}_{x}=\frac{2}{3}x+N_{\rho}^{(t+1)}(x)$, where $N_{\rho}^{(t+1)}(\cdot)$ accounts for any positive jumps due to $\rho^{(t+1)}$.
For each (fixed) $s$, we consider the stochastic process $G^{(t)}(s)$. To do this, one could analyze the two functions $H_{\beta}^{(t+1)}(s),\ H_{\alpha}^{(t+1)}(s)$ separately as follows. Notice that $\beta^{(t)}_{x}$ and $\alpha^{(t)}_{x}$ are (conditionally) mean-zero for each time $t$ and location $x$. The simpler term to analyze is $H_{\beta}^{(t+1)}(s)$. We have
$H_{\beta}^{(t+1)}(s)$ $=\int_{\mathds{R}}e^{-sx-s\beta^{(t+1)}_{x}}dF^{(t)}(x)=\int_{\mathds{R}}e^{-s\beta^{(t+1)}_{x}}e^{-sx}dF^{(t)}(x),$
and with a suitable MGF control, we expect that
$H_{\beta}^{(t+1)}(s)\leqslant C^{(t)}\cdot G^{(t)}(s),$
for some constants satisfying $\prod_{i=1}^{\infty}C^{(t)}<\infty$. Here, we use the fact that $\rho^{(t+1)}(x)$ is independent of the canonical filtration (up to time $t$). The term $H_{\alpha}^{(t+1)}(s)$ is more difficult to analyze due to the fact that the $\alpha^{(t+1)}_{\cdot}$ process is highly correlated with the $\beta^{(t+1)}_{\cdot}$ process and will be analyzed in a future paper.
4.1 Simulations of Conjecture 1
In Figure 5, we show simulations for two independent copies of our dynamics. We observe the conjectured convergence as $t$ increases. However, we obviously cannot pictorially depict the convergence of the measure over distribution functions.
(a) First Simulation
(b) Second Simulation
Figure 5: Empirical distribution of $\overleftarrow{F}^{(t)}$ across time for two different samples.
5 Future Work
In this paper, we analyze a simple dynamical clustering algorithm for $\mathds{R}$-valued data in which each data point attempts to cluster with either its left or right neighbor, chosen uniformly at random. For this model, we establish (under mild assumptions) a unique scaling limit which is independent of the initial data.
We conjecture that (under similar assumptions), a similar scaling limit holds for other realistic dynamic clustering algorithms, such as Algorithm 2 from the introduction of this paper. For this algorithm, the scaling limit appears to depend on the initial data. The proof technique used in this work does not apply to Algorithm 2, as the weight sequences in the time reversal are no longer described via integer weights. Determining the existence of stationary distributions for this type of dynamics remains an important area of future work. New techniques are required to study these algorithms in a greater generality.